# On rate-optimal partitioning classification from observable and from privatised data

### v2

@@ 3 Partitioning classification under local differential privacy
One of the main purposes of this paper is to bound the error probability of partitioning classifiers in the case, when the raw data $\mathcal{D}_{n}$ are not directly accessible, but only a suitably anonymised surrogate. More precisely, the anonymised data must satisfy a local differential privacy (LDP) condition [2, 5]. Our work is motivated by Berrett and Butucea’s recent paper [2], where the first step in this direction was done.
@@
$Z_{i,j}=Y_{i}{\mathbb{I}}_{\{X_{i}\in A_{h,j}\}}+\sigma_{Z}\epsilon_{i,j},\quad j=1,\ldots\vspace{0.5mm}$
@@
Now, we briefly recall the definition of LDP. A non-interactive privacy mechanisms can be described by the conditional distributions $Q_{i}$ of the privatized data $Z_{i}$, for $i=1,\dots,n$, where each $Z_{i}$ takes its values from a measurable space $(\mathcal{Z},\mathscr{Z})$. Specifically, given a realization of the raw data $(X_{i},Y_{i})=(x_{i},y_{i})$, one generates $Z_{i}$ according to the probability measure defined by $Q_{i}(A\mid(X_{i},Y_{i})=(x_{i},y_{i}))$, for any $A\in\mathscr{Z}$. Such a non-interactive mechanism is local since any data holder can independently generate privatized data (e.g., without a trusted third party). For a privacy parameter $\alpha\in[0,\infty]$, a non-interactive privacy mechanism is said to be an $\alpha$-locally differentially private mechanism if the condition
$\frac{Q_{i}(A\mid(X_{i},Y_{i})=(x,y))}{Q_{i}(A\mid(X_{i},Y_{i})=(x^{\prime},y^{\prime}))}\leq\exp(\alpha)\vspace{1mm}$
@@
The next theorem is the extension of Theorem 2.1 to the setup of privatised data.
@@
If $X$ is bounded, $m$ satisfies the Lipschitz condition (2), the margin condition (3) with $0\leq\gamma$ and the combined margin and density condition (9) such that $\sqrt{f_{h}}$ and $\sqrt{f}$ are replaced by $f_{h}$ and $f$, respectively, then
${\mathbb{E}}(L(\widetilde{D}_{n}))-L^{*}$ $=O\left(h^{1+\gamma}\right)+O\left(\left(\frac{\sigma_{Z}^{2}}{nh^{2d_{a}}}\right)^{(1+\min\{1,\gamma,\gamma(f)\})/2}\right)+O\left(\frac{\sigma_{Z}}{nh^{d}}\right).$
For the choice
$h=h_{n}=(n/\sigma_{Z}^{2})^{-\frac{1}{2+2d_{a}}},$
if $d\leq 2d_{a}$, one has
${\mathbb{E}}\{L(\widetilde{D}_{n})\}-L^{*}$ $=O\left((n/\sigma_{Z}^{2})^{-\frac{1+\min\{1,\gamma,\gamma(f)\}}{2+2d_{a}}}\right).$
For non-privatised data, $\gamma(\sqrt{f})$ is larger, than for privatised data $\gamma(f)$. Thus, it may happen that in a particular case for non-privatised data, $O\left(n^{-\frac{1+\gamma}{2+d_{a}}}\right)$ dominates $O\left(n^{-\frac{1+\gamma(\sqrt{f})}{2+d_{a}}}\right)$, while for privatised data, just the opposite, $O\left(n^{-\frac{1+\gamma(f)}{2+2d_{a}}}\right)$ dominates $O\left(n^{-\frac{1+\gamma}{2+2d_{a}}}\right)$.
@@
The next theorem is a multi-label extension of Theorem 3.1.
@@
If $X$ is bounded, $P_{1},\dots,P_{M}$ satisfy the Lipschitz condition (2), the margin condition (14) with $0\leq\gamma$ and the combined margin and density condition (15) such that $\sqrt{f_{h}}$ and $\sqrt{f}$ are replaced by $f_{h}$ and $f$, respectively, then
${\mathbb{E}}\{L(\widetilde{D}_{n})\}-L^{*}$ $=O\left(h^{1+\gamma}\right)+O\left(\Big{(}\frac{\sigma_{Z}^{2}}{nh^{2d_{a}}}\Big{)}^{(1+\min\{1,\gamma,\gamma(f)\})/2}\right)+O\left(\frac{\sigma_{Z}}{nh^{d}}\right).$
@@ 4 Proofs
cf. Theorem 2.2 in Devroye, Györfi and Lugosi [4]. For $x\in A_{h,j}$, we set
@@
and thus the bound on the approximation error in (11). Concerning the estimation error $J_{n}$, we have that
@@
where in the last step we refer to CLT. (At the end of this subsection we show that the error term for this CLT approximation is of order $O(1/(nh^{d_{a}}))$.)
@@
$e^{-t^{2}/2}\frac{1}{\sqrt{2\pi}}\frac{1}{t}\left(1-\frac{1}{t^{2}}\right)\leq\Phi(-t)\leq e^{-t^{2}/2}\frac{1}{\sqrt{2\pi}}\frac{1}{t},$
@@
for $h$ small enough. (We note that the CLT approximation is not needed for the discrete part.) One has that
@@
[…]
$=O\left(\frac{1}{(nh^{d_{a}})^{(\gamma+1)/2}}\right).$ (21)
(We note that the CLT approximation with an error term $O(1/(nh^{d_{a}}))$ is justified by (25) below.) For the notation (8) and for $a=nh^{d_{a}}/8$, the combined margin and density assumption implies that
@@
$=\frac{4}{\sqrt{nh^{d_{a}}}}\int_{0}^{\infty}e^{-as^{2}}G_{\sqrt{f_{h}}}(ds)$
[…]
$=O\left(\frac{1}{(nh^{d_{a}})^{(\gamma(\sqrt{f})+1)/2}}\right).$ (22)
($0<h\leq h_{0}$). (We note that the CLT approximation with an error term $O(1/(nh^{d_{a}}))$ is justified by (24) below.) (20), (4.1) and (22) together with (24) and (25) below yield the bound on the estimation error in (11). ∎
@@
For the CLT approximation in (4.1), we need an upper bound. Put $Z_{i,A}=Y_{i}{\mathbb{I}}_{\{X_{i}\in A\}}$. We use the Berry-Esséen theorem and the normal approximation to upper bound this probability. Recall that because of the nonuniform Berry-Esséen theorem there exists a universal constant $0.47850.4097<c<$ 0.4785 such that for all $a\in{\mathbb{R}}$ we have
@@
$\Delta$ $\leq O\left(\frac{1}{n}\right)\sum_{A\in{\cal P}_{h}}\frac{1}{\mu(A)}\mu(A)=O\left(\frac{1}{nh^{d}}\right).$
@@
$\Delta^{\prime}$ $=O\left(\frac{1}{n}\right)\sum_{A\in{\cal P}_{h}}\frac{1}{\mu(A)}\mu_{a}(A)\leq O\left(\frac{1}{n}\right)\sum_{A\in{\cal P}_{h}}\frac{1}{\mu_{a}(A)}\mu_{a}(A)=O\left(\frac{1}{nh^{d_{a}}}\right).$ (24)
@@
$\Delta^{\prime\prime}$ $=O\left(\frac{1}{n}\right)\sum_{A\in{\cal P}_{h}}\frac{1}{h^{d_{a}}}\mu(A)=O\left(\frac{1}{nh^{d_{a}}}\right).$ (25)
@@
$G_{\sqrt{f_{h}}}(t)$ $=\int_{S_{a}}{\mathbb{I}}_{\{0<{\sqrt{f_{h}(x)}|m(x)|\leq t\}}}\frac{1}{\sqrt{f_{h}(x)}}\mu_{a}(dx)$
@@
Put
$P_{n,k}(x)=\frac{\nu_{n,k}(A_{n,j})}{\mu(A_{n,j})},\quad\mbox{ if }x\in A_{n,j}.$
In the proof the main ingredient is Lemma 8 in [9]:
@@
Let $D_{n}$ be a plug-in rule with any estimates $P_{n,k}$ of $P_{k}$. Abbreviating
$R_{l}^{*}(x)=P_{D^{*}(x)}(x)-P_{l}(x)\geq 0,$
@@
We bound (26) by
@@
$J_{n,k,l}^{(1)}$ $=\int R_{l}^{*}(x){\mathbb{I}}_{\{l\neq g^{*}(x)\}}{\mathbb{P}}\{|P_{n,k}(x)-{\mathbb{E}}\{P_{n,k}(x)\}|\geq R_{l}^{*}(x)/(2M)\}\mu(dx),$
@@
$J_{n,k,l}^{(2)}$ $=\int R_{l}^{*}(x){\mathbb{I}}_{\{l\neq g^{*}(x)\}}{\mathbb{I}}_{\{|{\mathbb{E}}\{P_{n,k}(x)\}-P_{k}(x)|\geq R_{l}^{*}(x)/(2M)\}}\mu(dx).$
Concerning the estimation error $J_{n,k,l}^{(1)}$, as in the proof of Theorem 2.1 we apply the CLT with Berry-Esséen bound and then decomposition (7) and also (20). For $x\in A_{n,j}$, we have that
${\mathbb{P}}\{|P_{n,k}(x)-{\mathbb{E}}\{P_{n,k}(x)\}|\geq R_{l}^{*}(x)/(2M)\}$
=$={\mathbb{P}}\{|\nu_{n,k}(A_{n,j})-{\mathbb{E}}\{\nu_{n,k}(A_{h,j})\}|\geq\mu(A_{h,j})R_{l}^{*}(x)/(2M)\}$
$\approx 2\Phi\left(-\sqrt{n}\frac{\mu(A_{h,j})|R_{l}^{*}(x)|/(2M)}{\sqrt{{\mathbb{V}ar}(Y{\mathbb{I}}_{\{X\in A_{h,j}\}}})}\,\right),$
and henceforth, up to a term $O(1/(nh^{d_{a}}))$,
$\sum_{A\in{\cal P}_{h}}\int_{A}{\mathbb{P}}\Big{(}|\nu_{n,k}(A)-{\mathbb{E}}[\nu_{n,k}(A)]|\geq\mu(A)|R_{l}^{*}(x)|/(2M)\Big{)}|R_{l}^{*}(x)|\mu(dx)$
$\approx 2\sum_{A\in{\cal P}_{h}}\int_{A}\Phi\left(-\sqrt{n}\frac{\mu(A)|R_{l}^{*}(x)|/(2M)}{\sqrt{{\mathbb{V}ar}(Y{\mathbb{I}}_{\{X\in A\}}})}\,\right)|R_{l}^{*}(x)|\mu(dx)$
$\leq\frac{4M}{\sqrt{n}}\sum_{A\in{\cal P}_{h}}\int_{A}\exp\left(-\frac{n}{8M^{2}}\mu(A)R_{l}^{*}(x)^{2}\right)\frac{1}{\sqrt{\mu(A)}}\mu(dx)$
$\leq\frac{4M}{\sqrt{n}}\sum_{A\in{\cal P}_{h}}{\mathbb{I}}_{\{\mu(A)>h^{d_{a}}\}}\int_{A}\exp\left(-\frac{n}{8M^{2}}R_{l}^{*}(x)^{2}\mu(A)\right)\mu(dx)\frac{1}{\sqrt{\mu(A)}}$
$+\frac{4M}{\sqrt{n}}\sum_{A\in{\cal P}_{h}}{\mathbb{I}}_{\{h^{d_{a}}\geq\mu_{a}(A)>0\}}\int_{A}\exp\left(-\frac{n}{8M^{2}}R_{l}^{*}(x)^{2}\mu_{a}(A)\right)\mu_{a}(dx)\frac{1}{\sqrt{\mu(A)}}$
$+\frac{4M}{\sqrt{n}}\sum_{A\in{\cal P}_{h}}{\mathbb{I}}_{\{h^{d_{a}}\geq\mu_{s}(A)>0\}}\sqrt{\mu_{s}(A)},$
@@
$\frac{4M}{\sqrt{nh^{d_{a}}}}\int\exp\left(-\frac{nh^{d_{a}}}{8M^{2}}R_{l}^{*}(x)^{2}\right)\mu(dx)$
$+\frac{4M}{\sqrt{nh^{d_{a}}}}\int_{S_{a}}\exp\left(-\frac{nh^{d_{a}}}{8M^{2}}R_{l}^{*}(x)^{2}f_{h}(x)\right)\frac{1}{\sqrt{f_{h}(x)}}\mu_{a}(dx),$
for $h$ small enough. If $l\neq g^{*}(x)$, then $R_{l}^{*}(x)\geq P_{(1)}(x)-P_{(2)}(x)=:\Delta(x)$ otherwise $R_{l}^{*}(x)=0$, therefore
$J_{n,k,l}^{(1)}$ $\leq{\mathbb{I}}_{\{l\neq g^{*}(x)\}}\frac{4M}{\sqrt{nh^{d_{a}}}}\int\exp\left(-\frac{nh^{d_{a}}}{8M^{2}}\Delta(x)^{2}\right)\mu(dx)$
$+{\mathbb{I}}_{\{l\neq g^{*}(x)\}}\frac{4M}{\sqrt{nh^{d_{a}}}}\int_{S_{a}}\exp\left(-\frac{nh^{d_{a}}}{8M^{2}}\Delta(x)^{2}f_{h}(x)\right)\frac{1}{\sqrt{f_{h}(x)}}\mu_{a}(dx).$

### v3

@@ 3 Partitioning classification under local differential privacy
One of the main purposes of this paper is to bound the error probability of partitioning classifiers in the case, when the raw data $\mathcal{D}_{n}$ is not directly accessible, but only a suitably anonymised surrogate. More precisely, the anonymised data must satisfy a local differential privacy (LDP) condition [3, 7]. Our work is motivated by Berrett and Butucea’s recent paper [3], where the first step in this direction was done. Notice that the same privatization mechanism was studied for regression and density estimation in [12] and [11], respectively.
@@
$Z_{i,j}=Y_{i}{\mathbb{I}}_{\{X_{i}\in A_{h,j}\}}+\sigma_{Z}\epsilon_{i,j},\quad j=1,\ldots\vskip 1.42262pt$
@@
Now, we briefly recall the definition of LDP. Non-interactive privacy mechanisms can be described by the conditional distributions $Q_{i}$ of the privatized data $Z_{i}$, for $i=1,\dots,n$, where each $Z_{i}$ takes its values from a measurable space $(\mathcal{Z},\mathscr{Z})$. Specifically, given a realization of the raw data $(X_{i},Y_{i})=(x_{i},y_{i})$, one generates $Z_{i}$ according to the probability measure defined by $Q_{i}(A\mid(X_{i},Y_{i})=(x_{i},y_{i}))$, for any $A\in\mathscr{Z}$. Such a non-interactive mechanism is local since any data holder can independently generate privatized data (e.g., without a trusted third party). For a privacy parameter $\alpha\in[0,\infty]$, a non-interactive privacy mechanism is said to be an $\alpha$-locally differentially private mechanism if the condition
$\frac{Q_{i}(A\mid(X_{i},Y_{i})=(x,y))}{Q_{i}(A\mid(X_{i},Y_{i})=(x^{\prime},y^{\prime}))}\leq\exp(\alpha)\vskip 2.84526pt$
@@
Let us introduce the modified combined margin and density condition for privatization. We say that $m$ satisfies the modified combined margin and density condition if there exists $h_{2}^{*}$ such that for all $h\in(0,h_{2}^{*})$ we have for all $t>0$:
$\widetilde{G}_{h}(t)\doteq\int{\mathbb{I}}_{\{0<f_{h}(x)(P_{(1)}(x)-P_{(2)}(x))\leq t\}}\frac{1}{f_{h}(x)}\mu_{a}(dx)\leq c_{2}\cdot t^{\gamma_{2}}$ (16)
with $c_{2}>0$ and $\gamma_{2}\geq 0$.
The next theorem is the extension of Theorem 2.1 to the setup of privatised data. The rate of convergence differs in two ways w.r.t. the previous result. The estimation error grows with $1/h^{-2d_{a}}$, instead of $1/h^{-d_{a}}$ which was the case for (11). Besides, the rate of convergence depends on $\gamma_{2}$, which is the modified combined margin and density condition parameter. The proof of Theorem 3.1 is included in Section 5.
@@
If $X$ is bounded, $m$ satisfies the Lipschitz condition (2), the margin condition (3) with $\gamma\geq 0$ and the combined margin and density condition (9) with $\gamma_{1}$, the modified combined margin condition with $\gamma_{2}$. Let $(h_{n})$ be a monotonically decreasing sequence with zero limit, then we have
${\mathbb{E}}(L(\widetilde{D}_{n}))-L^{*}$ $=O\left(h_{n}^{1+\gamma}\right)+O\left(\bigg{(}\frac{1}{nh_{n}^{d_{a}}}\bigg{)}^{(\min\{1,\gamma,\gamma_{1})\}+1)/2}\right)$
$\quad+O\left(\left(\frac{\sigma_{Z}^{2}}{nh_{n}^{2d_{a}}}\right)^{(1+\min\{\gamma,\gamma_{2}\})/2}\right)+O\left(\frac{\sigma_{Z}}{nh_{n}^{d}}\right).$
If for all $n\in\mathbb{N}^{+}$:
$h_{n}\leq n^{\frac{\min(1,\gamma,\gamma_{1})-\min(\gamma,\gamma_{2})}{d_{a}(2\min(\gamma,\gamma_{2})+1-\min(1,\gamma,\gamma_{1}))}}$
and $d\leq 2d_{a}$, then the second and fourth terms are dominated by the third term, hence for
$h_{n}=(n/\sigma_{Z}^{2})^{-\frac{1}{2+2d_{a}}},$
one has
${\mathbb{E}}\{L(\widetilde{D}_{n})\}-L^{*}$ $=O\left((n/\sigma_{Z}^{2})^{-\frac{1+\min\{1,\gamma,\gamma_{2}\}}{2+2d_{a}}}\right).$
@@
Our final theorem, which is proved in Section 5, is the multi-label version of Theorem 3.1.
@@
If $X$ is bounded, $P_{1},\dots,P_{M}$ satisfy the Lipschitz condition (2), the margin condition (3) with $\gamma\geq 0$, the combined margin and density condition (9) with $\gamma_{1}$ and the modified combined margin condition with $\gamma_{2}$. Let $(h_{n})$ be a monotonically decreasing null sequence, then we have
${\mathbb{E}}\{L(\widetilde{D}_{n})\}-L^{*}$ $=O\left(h_{n}^{1+\gamma}\right)+O\left(\bigg{(}\frac{1}{nh_{n}^{d_{a}}}\bigg{)}^{(\min\{1,\gamma,\gamma_{1})\}+1)/2}\right)$
$\quad+O\left(\bigg{(}\frac{\sigma_{Z}^{2}}{nh_{n}^{2d_{a}}}\bigg{)}^{(1+\min\{\gamma,\gamma_{2}\})/2}\right)+O\left(\frac{\sigma_{Z}}{nh_{n}^{d}}\right).$
4 Discussion
In this paper we investigated both the binary and the multiclass versions of partitioning classification. Studying these methods also helps better understanding a wide range of local averaging estimators, and it is directly relevant for various statistical and machine learning methods, including federated learning and ensemble-based approaches. One of our main contributions was that we weakened the strong density assumption (SDA) and proved fast convergence rates under the margin condition and a newly introduced combined margin and density condition. It was shown that these convergence rates can be achieved even without using SDA, refuting the conjecture that without SDA the convergence rate can be arbitrarily slow.
We extended our results for the (locally differentially) private partitioning algorithm, which has no direct access to the data, and we proved convergence rates under the margin condition, combined margin and density condition and the modified combined margin and density condition. As expected, the convergence rates of the nonprivate algorithms are faster than the corresponding privatized ones. Our theorems quantify the effect of privatization by introducing an extra rate of $2$ for the bandwidth of the partitions and adjust the dependence of the rate to the density of the inputs by using a modified combined margin and density parameter.
Acknowledgements
The research of László Györfi was supported by the National Research, Development and Innovation Office (NKFIH) of Hungary under the 2023-1.1.1-PIACI-FÓKUSZ-2024-00051 funding scheme. The work of Balázs Csanád Csáji and Ambrus Tamás was supported, in part, by the European Union project RRF-2.3.1-21-2022-00004 within the framework of the Artificial Intelligence National Laboratory of Hungary.
@@ 5 Proofs
cf. Theorem 2.2 in Devroye, Györfi and Lugosi [6]. For notational simplicity let $h=h_{n}$. For $x\in A_{h,j}$, we set
@@
and thus the bound on the approximation error in (11).
Concerning the estimation error $J_{n}$, we have that
@@
where in the last step we refer to CLT. (At the end of this subsection we show that the error term for this CLT approximation is of order $O(1/(nh_{n}^{d_{a}}))$.)
@@
$\Phi(-t)\leq e^{-t^{2}/2}\frac{1}{\sqrt{2\pi}}\frac{1}{t},$
@@
for $h$ small enough. (We note that the CLT approximation is not needed for the discrete part.)
Additionally, one has that
@@
$=\frac{4}{\sqrt{nh^{d_{a}}}}2a\int_{0}^{\infty}se^{-as^{2}}G^{*}(s)ds$
$\leq\frac{4}{\sqrt{nh^{d_{a}}}}2c^{*}a\int_{0}^{\infty}e^{-as^{2}}s^{1+\gamma}ds$
$=\frac{4}{\sqrt{nh^{d_{a}}}}2c^{*}a^{-\gamma/2}\int_{0}^{\infty}e^{-u^{2}}u^{1+\gamma}du$
$=O\left(\frac{1}{(nh^{d_{a}})^{(\gamma+1)/2}}\right),$ (22)
where recall that $G^{*}(s)=\int{\mathbb{I}}_{\{0<|m(x)|\leq t\}}\mu(dx)$. (Note that the CLT approximation with an error term $O(1/(nh_{n}^{d_{a}}))$ is justified by (26) below.)
For notation (8) and for $a=nh^{d_{a}}/8$, the combined margin and density assumption implies that for $0<h\leq h_{1}^{*}$, we have
@@
$=\frac{4}{\sqrt{nh^{d_{a}}}}\int_{0}^{\infty}e^{-as^{2}}G_{h}(ds)$
$=\frac{4}{\sqrt{nh^{d_{a}}}}2a\int_{0}^{\infty}se^{-as^{2}}G_{h}(s)ds$
$\leq\frac{4}{\sqrt{nh^{d_{a}}}}2c_{1}a\int_{0}^{\infty}e^{-as^{2}}s^{1+\gamma_{1}}ds$
$=\frac{4}{\sqrt{nh^{d_{a}}}}2c_{1}a^{-\gamma_{1}/2}\int_{0}^{\infty}e^{-u^{2}}u^{1+\gamma_{1}}du$
$=O\left(\frac{1}{(nh_{n}^{d_{a}})^{(\gamma_{1}+1)/2}}\right),$ (23)
(We note that the CLT approximation with an error term $O(1/(nh_{n}^{d_{a}}))$ is justified by (25) below.) (21), (5.1) and (23) together with (25) and (26) below yield the bound on the estimation error in (11). ∎
@@
For the CLT approximation in (5.1), we need an upper bound. Put $Z_{i,A}=Y_{i}{\mathbb{I}}_{\{X_{i}\in A\}}$. We use the Berry-Esséen theorem and the normal approximation to upper bound this probability. Recall that because of the nonuniform Berry-Esséen theorem there exists a universal constant $0.4097<c<0.4785$ such that for all $a\in{\mathbb{R}}$ we have
@@
$\Delta$ $\leq O\left(\frac{1}{n}\right)\sum_{A\in{\cal P}_{h}}1=O\left(\frac{1}{nh_{n}^{d}}\right).$
@@
$\Delta^{\prime}$ $=O\left(\frac{1}{n}\right)\sum_{A\in{\cal P}_{h}}\frac{1}{\mu(A)}\mu_{a}(A)\leq O\left(\frac{1}{n}\right)\sum_{A\in{\cal P}_{h}}1=O\left(\frac{1}{nh_{n}^{d_{a}}}\right).$ (25)
@@
$\Delta^{\prime\prime}$ $=O\left(\frac{1}{n}\right)\sum_{A\in{\cal P}_{h}}\frac{1}{h^{d_{a}}}\mu(A)=O\left(\frac{1}{nh_{n}^{d_{a}}}\right).$ (26)
@@
$G_{h}(t)$ $=\int_{S_{a}}{\mathbb{I}}_{\{0<{\sqrt{f_{h}(x)}|m(x)|\leq t\}}}\frac{1}{\sqrt{f_{h}(x)}}\mu_{a}(dx)$
@@
Let
$P_{n,k}(x)=\frac{\nu_{n,k}(A_{h,j})}{\mu(A_{h,j})},\quad\mbox{ if }x\in A_{h,j}.$
The main ingredient of the proofs is the slight extension of Lemma 8 in [13]:
@@
Let $g_{n}$ be a plug-in rule with any estimates $P_{n,j}$ of $P_{j}$. For the notation
$\Delta_{l}(x)$ $=P_{(1)}(x)-P_{(l)}(x),$
@@
As in the proof of Lemma 8 in Györfi and Weiss [13],
${\mathbb{E}}\{L(g_{n})\}-L^{*}$
$=\int{\mathbb{E}}\{(P_{g^{*}(x)}(x)-P_{g_{n}(x)}(x)){\mathbb{I}}_{\{g^{*}(x)\neq g_{n}(x)\}}\}\mu(dx)$
$=\int{\mathbb{E}}\{(P_{g^{*}(x)}(x)-P_{g_{n}(x)}(x)){\mathbb{I}}_{\{P_{g^{*}(x)}(x)>P_{g_{n}(x)}(x)\}}{\mathbb{I}}_{\{P_{n,g_{n}(x)}(x)\geq P_{n,g^{*}(x)}(x)\}}\}\mu(dx).$
If $g^{*}(x)=j$ and $g_{n}(x)=l$, then
$\{P_{n,g_{n}(x)}(x)-P_{n,g^{*}(x)}(x)\geq 0\}$
$=\{P_{n,l}(x)-P_{l}(x)+P_{l}(x)-P_{g^{*}(x)}(x)+P_{j}(x)-P_{n,j}(x)\geq 0\}$
$\subset\left\{|P_{n,j}(x)-P_{j}(x)|\geq(P_{g^{*}(x)}(x)-P_{g_{n}(x)}(x))/2\right\}$
$\cup\left\{|P_{n,l}(x)-P_{l}(x)|\geq(P_{g^{*}(x)}(x)-P_{g_{n}(x)}(x))/2\right\}$
Therefore,
${\mathbb{E}}\{L(g_{n})\}-L^{*}$
$\leq\sum_{j=1}^{M}\int{\mathbb{E}}\{(P_{g^{*}(x)}(x)-P_{g_{n}(x)}(x)){\mathbb{I}}_{\{|P_{n,j}(x)-P_{j}(x)|\geq(P_{g^{*}(x)}(x)-P_{g_{n}(x)}(x))/2\}}\}\mu(dx)$
$\leq\sum_{j=1}^{M}\sum_{l=2}^{M}\int(P_{(1)}(x)-P_{(l)}(x)){\mathbb{E}}\{{\mathbb{I}}_{\{|P_{n,j}(x)-P_{j}(x)|\geq(P_{(1)}(x)-P_{(l)}(x))/2\}}\}\mu(dx)$
$=\sum_{j=1}^{M}\sum_{l=2}^{M}\int\Delta_{l}(x){\mathbb{P}}\{|P_{n,j}(x)-P_{j}(x)|\geq\Delta_{l}(x)/2\}\mu(dx).$
Proof.
As above let us use the simplified notation $h=h_{n}$. We bound (27) by
@@
$J_{n,k,l}^{(1)}$ $=\int\Delta_{l}(x){\mathbb{P}}\{|P_{n,k}(x)-{\mathbb{E}}\{P_{n,k}(x)\}|\geq\Delta_{l}(x)/4\}\mu(dx),$
@@
$J_{n,k,l}^{(2)}$ $=\int\Delta_{l}(x){\mathbb{I}}_{\{|{\mathbb{E}}\{P_{n,k}(x)\}-P_{k}(x)|\geq\Delta_{l}(x)/4\}}\mu(dx).$
Concerning the estimation error $J_{n,k,l}^{(1)}$, as in the proof of Theorem 2.1 we apply the CLT with the Berry-Esséen bound and then decomposition (7) and also (21). For $x\in A_{h,j}$, we have that
${\mathbb{P}}\{|P_{n,k}(x)-{\mathbb{E}}\{P_{n,k}(x)\}|\geq\Delta_{l}(x)/4\}$
$={\mathbb{P}}\{|\nu_{n,k}(A_{h,j})-{\mathbb{E}}\{\nu_{n,k}(A_{h,j})\}|\geq\mu(A_{h,j})\Delta_{l}(x)/4\}$
$\approx 2\Phi\left(-\sqrt{n}\frac{\mu(A_{h,j})|\Delta_{l}(x)|/4}{\sqrt{{\mathbb{V}ar}(Y{\mathbb{I}}_{\{X\in A_{h,j}\}}})}\,\right),$
and henceforth, up to a term $O(1/(nh_{n}^{d_{a}}))$,
$\sum_{A\in{\cal P}_{h}}\int_{A}{\mathbb{P}}\Big{(}|\nu_{n,k}(A)-{\mathbb{E}}[\nu_{n,k}(A)]|\geq\mu(A)|\Delta_{l}(x)|/4\Big{)}|\Delta_{l}(x)|\mu(dx)$
$\approx 2\sum_{A\in{\cal P}_{h}}\int_{A}\Phi\left(-\sqrt{n}\frac{\mu(A)|\Delta_{l}(x)|/4}{\sqrt{{\mathbb{V}ar}(Y{\mathbb{I}}_{\{X\in A\}}})}\,\right)|\Delta_{l}(x)|\mu(dx)$
$\leq\frac{8}{\sqrt{n}}\sum_{A\in{\cal P}_{h}}\int_{A}\exp\left(-\frac{n}{32}\mu(A)\Delta_{l}(x)^{2}\right)\frac{1}{\sqrt{\mu(A)}}\mu(dx)$
$\leq\frac{8}{\sqrt{n}}\sum_{A\in{\cal P}_{h}}{\mathbb{I}}_{\{\mu(A)>h^{d_{a}}\}}\int_{A}\exp\left(-\frac{n}{32}\Delta_{l}(x)^{2}\mu(A)\right)\mu(dx)\frac{1}{\sqrt{\mu(A)}}$
$\quad+\frac{8}{\sqrt{n}}\sum_{A\in{\cal P}_{h}}{\mathbb{I}}_{\{h^{d_{a}}\geq\mu_{a}(A)>0\}}\int_{A}\exp\left(-\frac{n}{32}\Delta_{l}(x)^{2}\mu_{a}(A)\right)\mu_{a}(dx)\frac{1}{\sqrt{\mu(A)}}$
$\quad+\frac{8}{\sqrt{n}}\sum_{A\in{\cal P}_{h}}{\mathbb{I}}_{\{h^{d_{a}}\geq\mu_{s}(A)>0\}}\sqrt{\mu_{s}(A)},$
@@
$\frac{8}{\sqrt{nh^{d_{a}}}}\int\exp\left(-\frac{nh^{d_{a}}}{32}\Delta_{l}(x)^{2}\right)\mu(dx)$
$+\frac{8}{\sqrt{nh^{d_{a}}}}\int_{S_{a}}\exp\left(-\frac{nh^{d_{a}}}{32}\Delta_{l}(x)^{2}f_{h}(x)\right)\frac{1}{\sqrt{f_{h}(x)}}\mu_{a}(dx),$
for $h$ small enough. If $l\neq g^{*}(x)$, then $\Delta_{l}(x)\geq P_{(1)}(x)-P_{(2)}(x)=:\Delta(x)$ otherwise $\Delta_{l}(x)=0$, therefore up to a Berry-Esséen error term we have
$J_{n,k,l}^{(1)}$ $\leq\frac{8}{\sqrt{nh^{d_{a}}}}\int\exp\left(-\frac{nh^{d_{a}}}{32}\Delta(x)^{2}\right)\mu(dx)$
$\quad+\frac{8}{\sqrt{nh^{d_{a}}}}\int_{S_{a}}\exp\left(-\frac{nh^{d_{a}}}{32}\Delta(x)^{2}f_{h}(x)\right)\frac{1}{\sqrt{f_{h}(x)}}\mu_{a}(dx).$
and also
$\sum_{k=1}^{M}\sum_{l=1}^{M}J_{n,k,l}^{(1)}$ $\leq\frac{8M^{2}}{\sqrt{nh^{d_{a}}}}\int\exp\left(-\frac{nh^{d_{a}}}{32}\Delta(x)^{2}\right)\mu(dx)$
$\quad+\frac{8M^{2}}{\sqrt{nh^{d_{a}}}}\int_{S_{a}}\exp\left(-\frac{nh^{d_{a}}}{32}\Delta(x)^{2}f_{h}(x)\right)\frac{1}{\sqrt{f_{h}(x)}}\mu_{a}(dx).$ (28)
Similarly to (5.1), the margin condition implies
$\frac{8M^{2}}{\sqrt{nh^{d_{a}}}}\int\exp\left(-\frac{nh^{d_{a}}}{32}(P_{(1)}(x)-P_{(2)}(x))^{2}\right)\mu(dx)$
$=\frac{8M^{2}}{\sqrt{nh^{d_{a}}}}\int_{0}^{\infty}e^{-as^{2}}G^{*}(ds)$
$=O\left((nh_{n}^{d_{a}})^{-(\gamma+1)/2}\right),$ (29)
where $a=(nh^{d_{a}})/32$. As (23), the combined margin and density condition yields that
$\frac{8M^{2}}{\sqrt{nh^{d_{a}}}}\int_{S_{a}}\exp\left(-\frac{nh^{d_{a}}}{32}(P_{(1)}(x)-P_{(2)}(x))^{2}f_{h}(x)\right)\frac{1}{\sqrt{f_{h}(x)}}\mu_{a}(dx)$
$=\frac{8M^{2}}{\sqrt{nh^{d_{a}}}}\int_{0}^{\infty}e^{-as^{2}}G_{h}(ds)$
$=O\left((nh_{n}^{d_{a}})^{-(\gamma_{1}+1)/2}\right).$ (30)